I have been trying to solve the following linear equation
3X-5Y=11
2x-6y=2
I keep getting the wrong answer so when I checked with you guys I noticed that it was substituted with 2(11-5y)/3-6y=2
I understand the substitution part but how do I solve from here what does the / mean next to the 3 what operation I'm suppose to there, I'm confused please help.

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

I have been trying to solve the following linear equation
3X-5Y=11
2x-6y=2 divide by 2 --> x - 3y = 1 --> x = 3y + 1
----
Sub for x in the 1st eqn
3X-5Y=11
3(3y+1) - 5y = 11
9y+3 - 5y = 11
4y = 8
y = 2
------
x = 3y + 1
x = 7

Answer by rothauserc(4718)

3x -5y = 11
2x -6y = 2
************************
solve second equation for x
add 6y to both sides of =
2x = 6y + 2
divide both sides of = by 2
x = 3y + 1
************************
now substitute for x in first equation
3(3y + 1) - 5y = 11
9y + 3 - 5y = 11
4y + 3 = 11
subtract 3 from both sides of =
4y = 8
divide both sides of = by 4
y = 2
******************************************
substitute for y in first equation
3x - 5(2) = 11
3x - 10 = 11
add 10 to both sides of =
3x = 21
divide both sides of = by 3
x = 7
*****************************************
solution is x = 7 and y = 2
****************************************
check this solution by substituting the solution in the second equation
2(7) -6(2) = 2
14 - 12 = 2
2 = 2
our solution checks